Two energy levels of an atomic system are separated by energy corresponding to frequency 5.0×1014 Hz. Assume that all atoms are in one or the other of these two energy levels. Compute the fraction of atoms in the upper energy level at temperature 600 K. Take kb = 1.38×10-23JK-1 and h=6.6×10-34Js.


1
Expert's answer
2021-06-15T14:18:15-0400

"\frac{n_1}{n_2}=\frac{h\nu}{kT}=\frac{6.63\cdot 10^{-34}\cdot 5\cdot 10^{14}}{1.38\cdot 10^{-23}\cdot 400}=0.4."